# Activation of entanglement in generalized entanglement swapping

### v1

@@ II.1.2 Unentangled measurements
It is not possible to find a measurement that can create non-zero entanglement in the $12|34$ bipartition of the post-measurement state, i.e., $\mathcal{C}_{n|12vs34}(|\Phi_{n}\rangle\langle\Phi_{n}|)>0$ (Eq.38), and $\mathcal{C}_{n|14vs23}(|\Phi_{n}\rangle\langle\Phi_{n}|)=0$ (Eq.32) for an element. This is because $\mathcal{C}_{n|14vs23}(|\Phi_{n}\rangle\langle\Phi_{n}|)=0$ implies that after the measurement, $\rho_{14|n}$ or $\rho_{23|n}$ has to be a pure state, indicating that the measurement element is either a projective or rank-1 POVM. This is due to the fact that $\mathcal{C}_{n|14vs23}=\sqrt{\frac{4}{3}[1-\mbox{Tr}(\rho_{14|n}^{2})]}=0$, which means $\mbox{Tr}(\rho_{14|n}^{2})=1$, establishing a direct relationship between the $\rho_{14|n}$ state and the measurement operator $\Pi_{n}$, as per Eq.6. If a measurement element is a rank-1 POVM or projective, then to obtain an unentangled measurement operator $\Pi_{n}$, it has to be in the form $|\tilde{i}\rangle\otimes|\tilde{j}\rangle$, where $|\tilde{i}\rangle$ and $|\tilde{j}\rangle$ are arbitrary bases. However, using this unentangled measurement operator form $|\tilde{i}\rangle\otimes|\tilde{j}\rangle$, one cannot create non-zero entanglement in the $12|34$ bipartition of $|\Phi_{n}\rangle$.
@@
If the first measurement $\mathcal{M}_{1}$ has non-zero entanglement in the $12|34$ bipartition of the post-measurement state (Eq. 5) for one outcome (i.e., $\mathcal{C}_{n|12vs34}(|\Phi_{n}\rangle\langle\Phi_{n}|)>0$ Eq. 38), and a second measurement $\mathcal{M}_{2}$ acts on the (2,3) pair again in succession, the final post-measurement state (Eq.29) between A and C may be entangled. We determine the entanglement criteria of the final post-measurement state $\rho_{14|nm}$ (Eq.29) between A and C in terms of measurement operators, which is described below.
To determine the entanglement of the final state $\rho_{14|nm}$ when Bob performs two arbitrary successive measurements $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ on qubits (2,3) of the initial state (Eq.1), we use the negativity measure to account for the arbitrariness of the measurements. We do not use the concurrence measure as it is difficult to obtain an expression of an arbitrary state without knowing more about the state’s parameters. Negativity of a bipartite state $\rho_{AB}$ is defined as
$\mathcal{N}(\rho_{AB})=\frac{\|\rho_{AB}^{T_{B}}\|-1}{2}$ (39)
Every orthonormal basis $|\phi_{n\alpha}\rangle$ and $|\psi_{m\delta}\rangle$ of different POVM elements can be written in the following computational basis form $|\phi_{n\alpha}\rangle=\sum_{i_{1},j_{1}=0}^{1}a_{n\alpha|i_{1}j_{1}}|i_{1}j_{1}\rangle$ and $|\psi_{m\delta}\rangle=\sum_{l_{1},k_{1}=0}^{1}b_{m\delta|l_{1}k_{1}}|l_{1}k_{1}\rangle$ where $\sum_{i_{1},j_{1}=0}^{1}|a_{n\alpha|i_{1}j_{1}}|^{2}=1$, $a_{n\alpha|i_{1}j_{1}}\in\mathbb{C}$ and $\sum_{l_{1},k_{1}=0}^{1}|b_{m\delta|l_{1}k_{1}}|^{2}=1$, $b_{m\delta|l_{1}k_{1}}\in\mathbb{C}$. The square root of any Hermitian, positive semidefinite matrix $U$ can be written as horn :
Let denote $U=(\rho_{14|nm}^{T_{B}})^{\dagger}\rho_{14|nm}^{T_{B}}$ (here, $T_{B}$ denote the partial transpose w.r.t. qubit 4 can be written as
$U$ $=$ $(\rho_{14|nm}^{T_{B}})^{\dagger}\rho_{14|nm}^{T_{B}}$ (47)
$\times$ $\mu_{m\delta}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\gamma|i_{5}j_{5}}b_{m\delta|k_{3}l_{3}}b^{*}_{m\delta|k_{4}l_{4}}a_{n\eta|i_{6}j_{6}}a_{n\beta|i_{3}j_{7}}a^{*}_{n\eta|i_{8}j_{4}}|i_{4}j_{3}\rangle\langle i_{8}j_{7}|$ (48)
where $p_{n}=\sum_{\alpha=1}^{4}\frac{\pi_{n\alpha}}{4}$ and $s_{nm}=\frac{1}{\sum_{\alpha=1}^{4}\pi_{n\alpha}}\sum_{\alpha,\delta=1}^{4}\sum_{i,j=0}^{1}\pi_{n\alpha}\mu_{m\delta}|a_{n\alpha|ij}b^{*}_{m\delta|ij}|^{2}$. We denote $X=\mbox{Tr}(U)$ and $Y=\det U$ which are given by:
$X$ $=$ $\mbox{Tr}(U)$ (55)
$\times$ $\mu_{m\delta}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\gamma|i_{5}j_{5}}b_{m\delta|k_{3}l_{3}}b^{*}_{m\delta|k_{4}l_{4}}a_{n\eta|i_{6}j_{6}}a_{n\beta|i_{3}j_{3}}a^{*}_{n\eta|i_{4}j_{4}}$ (56)
and
$Y$ $=$ $\det U$ (57)
$=$ $\sum_{e,f,g,h=1}^{4}\epsilon_{efgh}\tau_{1e}\tau_{2f}\tau_{3g}\tau_{4h}$
where $\epsilon_{efgh}$ is Levi-Civita symbol and $\tau_{1e}$, $\forall e\in\{1,2,3,4\}$ are the first row elements of $(\rho_{14|nm}^{T_{B}})^{\dagger}\rho_{14|nm}^{T_{B}}$ matrix. Similarly $\tau_{2f}$, $\tau_{3g}$, $\tau_{4h}$ $\forall e,f,g,h\in\{1,2,3,4\}$ are the 2nd, 3rd, 4th row elements of $(\rho_{14|nm}^{T_{B}})^{\dagger}\rho_{14|nm}^{T_{B}}$ matrix. Now putting $i_{4}=0$, $j_{3}=0$ in Eq.47 one can get the $\tau_{1e}$ and similarly, putting $i_{4}=0$, $j_{3}=1$, and $i_{4}=1$, $j_{3}=0$ and $i_{4}=1$, $j_{3}=1$ in Eq.47 one can find the $\tau_{2f}$, $\tau_{3g}$ and $\tau_{4h}$ expressions respectively (for detail calculation see Appendix B).
$\mathcal{N}(\rho_{14|nm})$ $=$ $\frac{\mbox{Tr}\sqrt{U}-1}{2}$ (58)
$=$ $\frac{X+4\sqrt{Y}}{2\sqrt{X+2\sqrt{Y}}}-\frac{1}{2}$
where $X$ and $Y$ are given above in Eq.55 and Eq.57 respectively. We get a relation between the entanglement of the post-measurement state and measurement parameters (Eq.58) of two successive measurements. When Bob performs an unentangled measurement, $\mathcal{M}_{1}$, which is unable to establish entanglement between the (1,4) pair, the entanglement in the $14|23$ bipartition is retained in the post-measurement state. A second measurement, $\mathcal{M}_{2}$, can then establish entanglement in the final post-measurement state, $\rho_{14|nm}$, if the negativity, $\mathcal{N}(\rho_{14|nm})$, is greater than zero.
@@
for $n=1,2,3,4$. The amount of entanglement of $\rho_{14|n}$ for $n\in\{1,2,3,4\}$ is $\mathcal{N}(\rho_{14|n})=\frac{3\lambda-1}{2}$. Note that, the state $\rho_{14|n}$ is separable for $0\leq\lambda\leq\frac{1}{3}$ and entangled for $\frac{1}{3}<\lambda\leq 1$. Hence, it is reflected that the state $\rho_{14|n}$ is entangled if and only if the measurement element is entangled. So, the average entanglement after completion of measurement $\mathcal{M}_{1}$ between A and C is
@@
Now using Eq.38, we obtain the entanglement in $12|34$ bipartition of post-measurement state $|\Phi_{n}\rangle$ (Eq.5) for each POVM elements which is given by
@@
In order to determine whether the second measurement $\mathcal{M}_{2}$ can disturb the post-measurement state $\rho_{14|n}$ of the first measurement $\mathcal{M}_{1}$ or not, we evaluate the $14|23$ bipartition entanglement of the post-measurement state of each element $\Pi_{n}$. Using Eq.32 we obtain the $14|23$ bipartition entanglement of the post-measurement state for each element $\Pi_{n}$ as:
@@
which is non-zero for $10\leq\lambda<$ 1 except at $\lambda=1$. When $\lambda=1$, the measurement becomes a perfect Bell basis measurement, which is a projective measurement. If a Bell measurement is performed initially on the (2,3) pair, no subsequent measurements can affect the post-measurement state $\rho_{14|n}$. However, when the measurement strength is between $10\leq\lambda<$ 1, a subsequent measurement can disturb the post-measurement state of the first measurement.
Now, after post-processing each measurement outcome of the first measurement (Eq.II.1.2), Bob performs another measurement $\mathcal{M}_{2}$ on the (2,3) pair. We consider the second measurement $\mathcal{M}_{2}$ (Eq.25) to be a purely separable operation. The POVM elements of $\mathcal{M}_{2}$ are defined as:
@@
Now, using Eq.39, we get the amount of entanglement of the state $\rho_{14|nm}$ which is given by
@@
For every outcome, the amount of entanglement is same. The average entanglement between A and C after the completion of two measurements $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ in succession is
@@
Figure 1: Variation of negativity ($\mathcal{N}_{avg}(\rho_{14|\mathcal{M}_{1}\mathcal{M}_{2}})$ and $\mathcal{N}_{avg}(\rho_{14|\mathcal{M}_{1}})$) with the measurement parameter $\lambda$ for the average entanglement between A and C. Here $\mathcal{N}_{avg}(\rho_{14|\mathcal{M}_{1}\mathcal{M}_{2}})$ represents the average entanglement established between A and C after the completion of two measurements $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ in succession. $\mathcal{N}_{avg}(\rho_{14|\mathcal{M}_{1}})$ represents the average entanglement between A and C after the completion of $\mathcal{M}_{1}$ measurement.
In Fig.1, we have plotted the average entanglement (Eq.67 and Eq.61) established between A and C with respect to the measurement parameter $\lambda$ after the completion of the successive two measurements $\mathcal{M}_{1}$ followed by $\mathcal{M}_{2}$ and after the completion of the first measurement $\mathcal{M}_{1}$ in terms of negativity measure. We make the following observations: (i) After the first measurement $\mathcal{M}_{1}$ the average entanglement $\mathcal{N}_{avg}(\rho_{14|\mathcal{M}_{1}})$ established between A and C is zero in between $0\leq\lambda\leq\frac{1}{3}$ cause the measurement is not entangled in that range. (ii) Due to two successive measurements the average entanglement $\mathcal{N}_{avg}(\rho_{14|\mathcal{M}_{1}\mathcal{M}_{2}})$ is non-zero for the whole range of the $0<\lambda\leq 1$ and increasing with the $\lambda$ where $\mathcal{M}_{2}$ is a separable operation. (iii) The average entanglement $\mathcal{N}_{avg}(\rho_{14|\mathcal{M}_{1}\mathcal{M}_{2}})$ established between A and C after the two successive measurements is greater than the first measurement for the whole range of the $\lambda$.
@@ II.2 Separable operations
A measurement that operates as a separable operation is incapable of transforming separable states into entangled ones rain1 . Therefore, if Bob initiates a measurement that constitutes a separable operation, it will not induce entanglement in the $12|34$ bipartition of the post-measurement state, signifying $\mathcal{C}_{n|12vs34}=0$ (Eq.38) for each POVM element. This implies that the post-measurement states of the (1,2) and (3,4) pairs must be pure for every outcome, as $\mathcal{C}_{n|12vs34}=0$ implies $\mbox{Tr}(\rho_{12|n}^{2})=1$ and $\mbox{Tr}(\rho_{34|n}^{2})=1$. Thus, in our scenario, the post-measurement states within the (1,2) and (3,4) pairs must remain pure.
@@
A separable operation $\{\Pi_{n}\}_{n=1}^{K}$ can be represented as:
@@
where $\sum_{n}^{K}\Pi_{n}=\mathbb{I}$ and $\mathcal{A}_{n}$, $\mathcal{B}_{n}$ are positive operators acting on $\mathcal{H}_{\mathcal{A}}$, $\mathcal{H}_{\mathcal{B}}$ Hilbert spaces respectively. In our scenario, if Bob conducts a separable operation on the (2,3) pair and the operators $\mathcal{A}_{n}$ and $\mathcal{B}_{n}$ are rank-1 for each element, there will be no second measurement $\mathcal{M}_{2}$ capable of disturbing the (1,4) pair. This is because a rank-one measurement on each qubit destroys the initial entanglement of $|\phi^{+}\rangle_{12}$ and $|\phi^{+}\rangle_{34}$ states.
@@
Suppose a single-qubit measurement element, $\mathcal{A}_{n}$, acts on qubit 2 of the initial state $|\Psi\rangle_{12}=\sum_{i,j=0}^{1}a_{ij}|ij\rangle$, where $\sum_{ij=0}^{1}|a_{ij}|^{2}=1$. In our case, $|\Psi\rangle_{12}=|\phi^{+}\rangle_{12}=\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle)$. Similarly, $\mathcal{B}_{n}$ acts on qubit 3 of the initial state $|\Psi\rangle_{34}=|\phi^{+}\rangle_{34}$. We assess the post-measurement state for $\mathcal{A}_{n}$ while retaining the arbitrary nature of the measurement. The same methodology applies to $\mathcal{B}_{n}$.
@@
From Eq.69, it’s evident that when $\alpha=1$, indicating a projective measurement or a rank-1 POVM, the post-measurement state is always separable. However, when $\alpha$ is greater than 1, there’s a possibility that the post-measurement state remains entangled. By considering arbitrary orthonormal bases for each element: $|\varphi_{n1}\rangle=\cos\frac{\theta_{n}}{2}|0\rangle+e^{i\phi_{n}}\sin\frac{\theta_{n}}{2}|1\rangle$ and $|\varphi_{n2}\rangle=\sin\frac{\theta_{n}}{2}|0\rangle-e^{i\phi_{n}}\cos\frac{\theta_{n}}{2}|1\rangle$ on $\mathbb{C}^{2}$, we evaluate the I-concurrence rungta of the post-measurement state of each outcome using Eq.30:
@@
Our protocol is also applicable when starting with two copies of $d\times d$ maximally entangled states, instead of two-qubit maximally entangled states. S. Yokoyama et al. yokoyama have shown that by considering the $d\times d$ maximally entangled states, the post-measurement state in the (1,4) pair will be $\frac{\Pi_{n}^{*}}{\mbox{Tr}(\Pi_{n})}$ if a generalized POVM element $\Pi_{n}$ is done on the (2,3) pair. The proof can be easily carried out by following our results and considering that every POVM element can be decomposed into $d^{2}$ complete orthonormal basis on $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ and $\pi_{n\alpha}\geq 0$ for all $n\in{1,2,...,K}$. To calculate $\mathcal{C}_{n|14vs23}$ for each POVM element using Eq.32, one must follow the same procedure, but with the sum running from $\alpha=1$ to $\alpha=d^{2}$ for $d\times d$ maximally entangled states. Similarly, in the case of Eq.38, one must sum for $i,j,l,k,i^{\prime},k^{\prime}=0,...,d-1$ and for $\alpha,\beta,\gamma,\eta=1,...,d^{2}$. Here, Lemma-1 and Lemma-2 are also applicable in this scenario. Therefore, our protocol is also valid in this scenario, if one starts with $d\times d$ maximally entangled states. However, in this scenario, one must deal with bound entangled states and PPT measurements horodecki1, among other considerations.

### v2

@@ II.1.2 Unentangled measurements
It is not possible to find a measurement that can create non-zero entanglement in the $12|34$ bipartition of the post-measurement state, i.e., $\mathcal{C}_{n|12vs34}(|\Phi_{n}\rangle\langle\Phi_{n}|)>0$, and $\mathcal{C}_{n|14vs23}(|\Phi_{n}\rangle\langle\Phi_{n}|)=0$ for an element. $\mathcal{C}_{n|14vs23}=\sqrt{\frac{4}{3}[1-\mbox{Tr}(\rho_{14|n}^{2})]}=0$ implies that $\mbox{Tr}(\rho_{14|n}^{2})=1$, indicating that the post-measurement states $\rho_{14|n}$ must be pure, suggesting the presence of a projective or rank-1 POVM element. For an unentangled rank-1 POVM element or an unentangled projective element, $\Pi_{n}$ must be in the form of a tensor product, thereby preventing the generation of non-zero entanglement in the $12|34$ bipartition of $|\Phi_{n}\rangle$.
@@
If the first measurement $\mathcal{M}_{1}$ results in non-zero entanglement in the $12|34$ bipartition of the post-measurement state (Eq. 5) for one outcome (i.e., $\mathcal{C}_{n|12vs34}(|\Phi_{n}\rangle\langle\Phi_{n}|)>0$), and a second measurement $\mathcal{M}_{2}$ acts on the (2,3) pair successively, the final post-measurement state (Eq.11) between A and C will become entangled with a suitable choice of measurements. We determine the entanglement criteria of the final post-measurement state $\rho_{14|nm}$ (Eq.11) between A and C in terms of measurement operators. We derive the exact expression of the negativity vidal of the post-measurement state of the (1,4) pair after the sequential measurements in Appendix C. To evaluate the entanglement of the final state $\rho_{14|nm}$, we use the negativity measure vidal to accommodate the arbitrariness of the measurements. We do not utilize the concurrence measure due to the difficulty in obtaining an expression for an arbitrary state without further information about the state’s parameters.
When Bob performs an unentangled measurement, $\mathcal{M}_{1}$, which is unable to establish entanglement between the (1,4) pair, the entanglement in the $14|23$ bipartition is retained in the post-measurement state. A second measurement, $\mathcal{M}_{2}$, can then establish entanglement in the final post-measurement state, $\rho_{14|nm}$, if the negativity, $\mathcal{N}(\rho_{14|nm})$, is greater than zero.
@@
for $n=1,2,3,4$. The amount of negativity of $\rho_{14|n}$ for $n\in\{1,2,3,4\}$ is
$\mathcal{N}(\rho_{14|n})=\frac{3\lambda-1}{2}.$ (20)
Note that, the state $\rho_{14|n}$ is separable for $0\leq\lambda\leq\frac{1}{3}$ and entangled for $\frac{1}{3}<\lambda\leq 1$. Hence, it is reflected that the state $\rho_{14|n}$ is entangled if and only if the measurement element is entangled. So, the average negativity after completion of measurement $\mathcal{M}_{1}$ between A and C is
@@
Now using Eq.17, we obtain the concurrence of $12|34$ bipartition of post-measurement state $|\Phi_{n}\rangle$ (Eq.5) for each POVM elements which is given by
@@
In order to determine whether the second measurement $\mathcal{M}_{2}$ can disturb the post-measurement state $\rho_{14|n}$ of the first measurement $\mathcal{M}_{1}$ or not, we evaluate the concurrence of $14|23$ bipartition of the post-measurement state of each element $\Pi_{n}$. Using Eq.14 we obtain the concurrence of $14|23$ bipartition entanglement of the post-measurement state for each element $\Pi_{n}$ as:
@@
which is non-zero for $10\leq\lambda<$ 1 except at $\lambda=1$. When $\lambda=1$, the measurement becomes a perfect Bell basis measurement, which is a projective measurement. If a Bell measurement is performed initially on the (2,3) pair, no subsequent measurements can affect the post-measurement state $\rho_{14|n}$. However, when the measurement parameter $\lambda$ is between $10\leq\lambda<$ 1, a subsequent measurement can disturb the post-measurement state of the first measurement.
Now, after post-processing each measurement outcome of the first measurement (Eq.II.1.2), Bob performs another measurement $\mathcal{M}_{2}$ on the (2,3) pair. We consider the second measurement $\mathcal{M}_{2}$ to be a purely separable operation. The POVM elements of $\mathcal{M}_{2}$ are defined as:
@@
We get the amount of negativity of the state $\rho_{14|nm}$ which is given by
@@
For every outcome, the amount of negativity is same. The average negativity between A and C after the completion of two measurements $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ in succession is
@@
Figure 1: Variation of negativity with the measurement parameter $\lambda$ for the average entanglement between A and C. Here $\mathcal{N}_{avg}(\rho_{14|\mathcal{M}_{1}\mathcal{M}_{2}})$ represents the average negativity established between A and C after the completion of two measurements $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ in succession. $\mathcal{N}_{avg}(\rho_{14|\mathcal{M}_{1}})$ represents the average negativity between A and C after the completion of $\mathcal{M}_{1}$ measurement.
In Fig.1, we have plotted the average entanglement (Eq.27 and Eq.21) established between A and C with respect to the measurement parameter $\lambda$ after the completion of the successive two measurements $\mathcal{M}_{1}$ followed by $\mathcal{M}_{2}$ and after the completion of the first measurement $\mathcal{M}_{1}$ in terms of negativity measure. We make the following observations: (i) After the first measurement $\mathcal{M}_{1}$ the average negativity $\mathcal{N}_{avg}(\rho_{14|\mathcal{M}_{1}})$ established between A and C is zero in between $0\leq\lambda\leq\frac{1}{3}$ because the measurement is not entangled in that range. (ii) Due to two successive measurements the average negativity $\mathcal{N}_{avg}(\rho_{14|\mathcal{M}_{1}\mathcal{M}_{2}})$ is non-zero for the whole range of the $0<\lambda\leq 1$ and increasing with the $\lambda$ where $\mathcal{M}_{2}$ is a separable operation. (iii) The average negativity $\mathcal{N}_{avg}(\rho_{14|\mathcal{M}_{1}\mathcal{M}_{2}})$ established between A and C after the two successive measurements is greater than the first measurement for the whole range of the $\lambda$.
@@ II.2 Separable operations
A measurement that operates as a separable operation is incapable of transforming separable states into entangled ones rain1 . Therefore, if Bob initiates a measurement that constitutes a separable operation, it will not induce entanglement in the $12|34$ bipartition of the post-measurement state, signifying $\mathcal{C}_{n|12vs34}=0$ (Eq.17) for each POVM element. This implies that the post-measurement states of the (1,2) and (3,4) pairs must be pure for every outcome, as $\mathcal{C}_{n|12vs34}=0$ implies $\mbox{Tr}(\rho_{12|n}^{2})=1$ and $\mbox{Tr}(\rho_{34|n}^{2})=1$. Thus, in our scenario, the post-measurement states within the (1,2) and (3,4) pairs remain pure.
@@
A separable operation $\{\Pi_{n}\}$ can be represented as:
@@
where $\sum_{n}\Pi_{n}=\mathbb{I}$ and $\mathcal{A}_{n}$, $\mathcal{B}_{n}$ are positive operators acting on $\mathcal{H}_{\mathcal{A}}$, $\mathcal{H}_{\mathcal{B}}$ Hilbert spaces respectively. In our scenario, if Bob conducts a separable operation on the (2,3) pair and the operators $\mathcal{A}_{n}$ and $\mathcal{B}_{n}$ are rank-1 for each element, there will be no second measurement $\mathcal{M}_{2}$ capable of disturbing the (1,4) pair. This is because a rank-one element on each qubit destroys the initial entanglement of $|\phi^{+}\rangle_{12}$ and $|\phi^{+}\rangle_{34}$ states.
@@
Suppose a single-qubit measurement element, $\mathcal{A}_{n}$, acts on qubit 2 of the initial state $|\Psi\rangle_{12}=|\phi^{+}\rangle_{12}=\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle)$. Similarly, $\mathcal{B}_{n}$ acts on qubit 3 of the initial state $|\Psi\rangle_{34}=|\phi^{+}\rangle_{34}$. We assess the post-measurement state for $\mathcal{A}_{n}$ while retaining the arbitrary nature of the measurement. The same methodology applies to $\mathcal{B}_{n}$ also.
@@
From Eq.29, it is evident that when the measurement element is a rank-1 POVM or projective, the post-measurement state is always separable. However, when the measurement element’s rank is greater than one, there is a possibility that the post-measurement state remains entangled. By considering arbitrary orthonormal bases for each element: $|\varphi_{n1}\rangle=\cos\frac{\theta_{n}}{2}|0\rangle+e^{i\phi_{n}}\sin\frac{\theta_{n}}{2}|1\rangle$ and $|\varphi_{n2}\rangle=\sin\frac{\theta_{n}}{2}|0\rangle-e^{i\phi_{n}}\cos\frac{\theta_{n}}{2}|1\rangle$ on $\mathbb{C}^{2}$, we evaluate the I-concurrence rungta of the post-measurement state of each outcome using Eq.12:
@@
Our protocol is also applicable when starting with two copies of $d\times d$ maximally entangled states, instead of two-qubit maximally entangled states. S. Yokoyama et al. yokoyama have shown that by considering the $d\times d$ maximally entangled states, the post-measurement state in the (1,4) pair will be $\frac{\Pi_{n}^{*}}{\mbox{Tr}(\Pi_{n})}$ if a generalized POVM element $\Pi_{n}$ is done on the (2,3) pair. The proof can be easily carried out by following our results and considering that every POVM element can be decomposed into $d^{2}$ complete orthonormal basis on $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ and $\pi_{n\alpha}\geq 0$ for all $n$. To calculate $\mathcal{C}_{n|14vs23}$ for each POVM element using Eq.14, one must follow the same procedure, but with the sum running from $\alpha=1$ to $\alpha=d^{2}$ for $d\times d$ maximally entangled states. Similarly, in the case of Eq.17, one must sum for $i,j,l,k,i^{\prime},k^{\prime}=0,...,d-1$ and for $\alpha,\beta,\gamma,\eta=1,...,d^{2}$. Here, Lemma-1 and Lemma-2 are also applicable in this scenario. Therefore, our protocol is also valid in this scenario, if one starts with $d\times d$ maximally entangled states. However, in this scenario, one must deal with bound entangled states and Positive Partial Transpose (PPT) measurements virmani ; horodecki1 ; horodecki2 ; bennett1 , among other considerations.